OpenStudy (anonymous):

. Evaluate: t^-ca^ca, where c = 3, a = 2, and t = -4

OpenStudy (anonymous):

just substitute \[\LARGE t^{-c}\cdot a^c\cdot a\] \[\LARGE (-4)^{-3}\cdot 2^3\cdot 2\] \[\LARGE \left(\frac{1}{-4}\right)^{3}\cdot 16\] Continue... :)

OpenStudy (anonymous):

the ^c is suppose to be with the (a) that you multiplied with at the end

OpenStudy (anonymous):

well.. I'm confused, O_O you should have used parentheses ... is it \[\Huge t^{-ca^{ca}}\] can you draw it ? :(

OpenStudy (anonymous):



OpenStudy (anonymous):

ok...